SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (24th April, 2007)

Attendees: Rainer Bartoldus, Al Eisner, Selina Li, Su Dong, Mark Tibbetts.

Operations

(From the note Debbie left for us)
  • JAS was updated to include new instructions for GLT transmission errors and GLT-FCT word corruption, asking the shifter to check the #events damged compared to #events seen by LiveFastMon and making the decision to contact experts accordingly.
  • Jim said in yesterday's 3.45 meeting that he's found a fix for this DHP disconnect problem that the L3 servers were seeing. It's not yet implemented.
  • Debbie is away until the 3rd, and she had changed the 'contact experts' page accordingly. Selina is taking the pager from now till Thursday and Hojeong will take it from then till Debbie is back.

EMT

  • The two broken TPB boards are fixed. Mark mentioned that we would like to test them at IR2 during a ROD. The chips on the board now stand higher than before, so it could be knocked off easily in the future if not handle properly. We need to make sure that the people who replace these TPB boards are aware of this. Su Dong suggested that we only use these boards at the teststand.
  • Mark will return to UK for about one month to renew his visa (May 4th - June 3rd). We need someone else to take the EMT cell phone during this period.

GLT

  • GLT lib validation is underway.
  • Last Wednesday after Debbie gave the monthly trigger status report at IR2, Bill and Walt suggested not to back off the delay to get rid of the low rate corruptions in the GLT diagnostic data. So, no action will be taken.

Trigger Configuration

  • SP9 data is appearing and needs validation. We will be able to use these MC samples to study the Zk trigger.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.